# Harmonic Mobile Manipulation

### v1

@@ Abstract
Recent advancements in robotics have enabled robots to navigate complex scenes or manipulate diverse objects independently. However, robots are still impotent in many household tasks requiring coordinated behaviors such as opening doors. The factorization of navigation and manipulation, while effective for some tasks, fails in scenarios requiring coordinated actions. To address this challenge, we introduce, HarmonicMM, an end-to-end learning method that optimizes both navigation and manipulation, showing notable improvement over existing techniques in everyday tasks. This approach is validated in simulated and real-world environments and adapts to novel unseen settings without additional tuning. Our contributions include a new benchmark for mobile manipulation and the successful deployment in a real unseen apartment, demonstrating the potential for practical indoor robot deployment in daily life.
Figure 1: Harmonic Mobile Manipulation: In this work, we address diverse mobile manipulation tasks integral to human’s daily life. Trained in a photo-realistic simulation, Our controller effectively accomplishes these tasks through harmonious mobile manipulation techniques in a real-world apartment featuring a novel layout, without any fine-tuning or adaptation.
00footnotetext: Work done during internship at PRIOR @ Allen Institute for AI00footnotetext: More results are on our project site: https://rchalyang.github.io/HarmonicMM
1 Introduction
The field of robot learning has traditionally treated robot navigation and manipulation as separate domains, leading to significant advancements in each. On one hand, agents have been trained to efficiently explore and navigate environments [19, 6, 11, 47], while on the other, substantial progress has been made in performing complex manipulation tasks, such as handling articulated objects with static arms in tabletop settings [26, 40, 51, 32].
However, in household settings, many tasks require coordinated body movements while simultaneously manipulating objects with our arms. To address this challenge, recent efforts have focused on enhancing robotic agents with the ability for mobile manipulation. The most prominent approach is invoking navigation and manipulation as separate modules. This separation is achieved either by utilizing a high-level semantic controller, such as a large language model [1, 16], or by estimating low-level subgoals using motion planners [44]. In these works, behaviors are split into distinct navigation and manipulation skills.
@@ 1 Introduction
  • Successful transfer of agents trained in simulation to real-world settings.
@@
2 Related Work
@@
Embodied AI Benchmark. Over the past few years, a variety of standard benchmarks have emerged to assess progress in the field of embodied AI. These benchmarks primarily focus on high-level planning and scene understanding skills of agents, encompassing tasks such as object navigation [20, 5, 46, 45, 33], point navigation, and house rearrangement [48, 31, 7, 9]. In these benchmarks, the interaction with the physical world is generally restricted to basic pick-and-place operations. Deitke et al.[5] have created diverse houses for training embodied agents to navigate through various everyday scenes. Yenamandra et al. [48] evaluate open vocabulary mobile manipulation, particularly focusing on the picking and placing of objects. Li et al.[23] tackles various household tasks using predefined motion primitives. In contrast to these benchmarks, our work extends beyond simple pick-and-place tasks. Our robot is designed to perform complex mobile manipulations, requiring tight coordination between navigation and manipulation.
Robotics Manipulation. In the field of robotic manipulation, research has predominantly focused on two areas: tabletop manipulation with a fixed robot base [50, 14, 28, 4, 3, 27] and simple mobile manipulation within constrained environments [28, 39, 2, 30]. Specifically, Mu et al.[28] have developed benchmarks for straightforward mobile manipulation tasks in open spaces. Gu et al.[13] have demonstrated how robots can clean tables using imitation learning in a tabletop setup. Urakami et al.[39] have created benchmarks for door-pulling tasks in restricted laboratory settings. Recent studies [2, 30] have shown how mobile robots can operate small closet doors by mimicking human actions. Our work, however, goes beyond the constraints of simple tabletop scenarios or limited base movement. We address more complex tasks that require active navigation within the environment, such as approaching and manipulating doors while in motion.
@@
Mobile Manipulator. In the realm of mobile manipulation, previous studies have designed task-specific mobile manipulators using optimization techniques[21] or employing Task-and-motion-planning (TAMP) methods [35, 10]. For instance, Lew et al.[21] developed a table-cleaning controller using whole-body trajectory optimization. While effective for their intended purposes, these methods tend to be narrowly focused with limited scalability and adaptability to different scenarios. In contrast, our pipeline could apply to a wide range of tasks, beyond the constraints of task-specific methodologies and offer a more flexible solution for various mobile manipulation challenges.
@@
Another significant line of work involves learning-based methods[44, 12], which offer better generalization in task completion. These methods diverge into two sub-categories: short-horizon whole-body control of robots [36, 15, 8], and long-horizon tasks solved at the scene level with predefined motion primitives in an iterative or two-stage manner [49, 44, 22, 17, 1, 4, 3]. Hu et al.[15] focus on tracking a given target pose with a whole-body mobile manipulator. Yokoyama et al.[49], Xia et al.[44] and Ahn et al.[1] exemplify solving long-horizon mobile manipulation tasks using predefined navigation and manipulation skills, either iteratively or in a two-stage approach. However, many everyday tasks defy the easy definition of a set of primitives and require close coordination between navigation and manipulation skills. Tasks such as opening doors and cleaning tables cannot be readily solved with predefined primitives. To address these challenges, our work proposes to jointly learn navigation and manipulation skills in daily scenes to solve complex mobile manipulation tasks that require close navigation and manipulation coordination.
3 Daily Mobile Manipulation Task Suite
We define Daily Mobile Manipulation Task Suite as the family of long-horizon tasks that necessitate close coordination between navigating one’s body and manipulating an object within the environment. Humans perform such tasks daily – from emptying a dishwasher and vacuuming an apartment to operating a vehicle and playing a racket sport. While we can effortlessly coordinate actions between our hands and feet, teaching robots to do the same is a very challenging problem. To study this problem, we present a suite of tasks, named the Daily Mobile Manipulation Task Suite, comprising atomic tasks that one might want a household robot to perform, as in Fig. 1. These tasks can neither be performed by a navigator with no arm nor by a manipulator with a fixed base. Instead, they require the agent to navigate its world while manipulating an object.
Task 1: Opening A Door. In most living spaces, navigating through doors is a key task. This involves identifying and operating the door handle, either by pulling or pushing, while adjusting body position to avoid collision with the door. The suite includes both Opening Door (Push) and Opening Door (Pull).
Task 2: Cleaning A Table. A common chore is cleaning surfaces like tables and counters requiring coordinated movement to clean the entire surface area effectively. The suite features Cleaning Table to represent such activities.
Task 3: Opening A Fridge. Operating household appliances is crucial for a robotic assistant. This task involves locating and opening an appliance like a fridge, including navigation and manipulation. The suite includes Opening Fridge as an example of such tasks.
Training Environments.We train agents in the Procedural AI2-THOR simulated environment to perform tasks within the Daily Mobile Manipulation Task Suite and evaluate these learned policies in the real world. To maximize generalization and reduce the simulation-to-real gap, we use a large collection of simulation worlds with photo-realistic visual observations. Specifically, we train our robot to open the door in 700 distinct houses, each with two rooms connected by a single openable door. For cleaning tables, we use 2000 houses with one to three rooms and a single table to clean. For opening a fridge, we use 1400 houses, each with a realistic kitchen populated with appliances, crockery, cutlery, and food items. We also randomize the texture of the floor and ceiling of the environment, as well as the objects, and vary the lighting in these simulated worlds. This high degree of variability enables our robot to generalize from simulated worlds to the real world.
4 Harmonic Mobile Manipulation (HarmonicMM)
Figure 2: Action Space Visualization: Our controller fully control all degree of freedom of the robot at every time-step.
@@ 4 Harmonic Mobile Manipulation (HarmonicMM)
Figure 3: HarmonicMM Network Architecture: Our HarmonicMM controller takes robot proprioception and multi-view visual observations as input and output navigation and manipulation commands at the same time. Every visual observation is preprocessed by a frozen DINOv2 encoder followed by a separate small CNN, and fed into the GRU and policy along with the proprioception reading.
Coordinating navigation and manipulation actions is challenging for embodied agents. Training agents with reinforcement learning has proven challenging with large action spaces. To simplify the problem, many previous works [1, 3, 48] perform tasks in a two-stage manner, where the robot navigates to a specific location and then manipulates objects with its base fixed. Some work [44, 36, 43] allow switching between the navigation and manipulation phases in an iterative manner; however, at each time step, only a restricted action space is invoked by the agent. This factorization of the action space prevents the robot from performing challenging tasks such as those in Daily Mobile Manipulation Task Suite.
We use the Stretch mobile manipulator [18] which is capable of navigating through scenes and manipulating objects simultaneously. As shown in Fig. 2, action $a\in\mathcal{R}^{5}$ draws from a 5-dimensional continuous action space, allowing us to command the Stretch robot to move its base and arm at each time step. Two navigation dimensions command the robot to translate its base for a specified distance, either forward or backward, and to rotate the base by a given angle, either clockwise or counterclockwise. Two manipulation dimensions control the arm’s height, up and down, as well as the arm’s extension, in and out. A fifth dimension controls the rotation of the wrist. For the Opening A Door and Opening A Fridge tasks, an additional sixth dimension corresponds to the invocation of the grasping action, which enables the robot to grasp the door knob or the edge of the fridge to interact with the given object.
4.1 Simulated Environment.
@@
While simulated environments like IsaacGym  [25] and Mujoco [38] support articulated manipulation and table cleaning, to the best of our knowledge, they lack the support for visual diversity that is essential for transferring skills to the real world for long-horizon planning. To enable the robot to learn the diverse tasks described in Sec 3 using ProcThor [5] for its visual diversity, we augmented the simulation environment with procedural door/fridge opening logic and photo-realistic table cleaning functionality.
@@
Procedural Door/Fridge Opening: In the ProcThor[5] simulation environment, doors and fridges are treated as semantic objects, and their open and closed states are not governed by any physical rules. Instead, we estimate their status at each simulation step based on the robot’s actions. For pushing actions, the system calculates the direction the robot’s end-effector is moving towards. If this direction aligns with the door’s normal vector, the door is opened incrementally in the same direction as the end-effector’s movement. In the case of pulling actions, the door’s movement is determined after the robot command is executed. The door opens in small segments if the end-effector’s moved direction aligns with the door’s normal vector, provided the robot has already secured the doorknob with a magnetic grasper. Additionally, collision detection algorithms are in place to prevent the door from clipping through the robot during these interactions.
@@ 4.2 Robot Perception
Navigating around a scene, interacting with different objects within the environment, and avoiding obstacles and collisions, require perceiving the world and being aware of one’s body. We, therefore, equip the robot with both RGB visual observations as well as the proprioceptive state of the robot, which is available on most robots and easy to obtain.
Multi-camera RGB Observations: Past works in mobile manipulation have employed various sensors, primarily dictated by the sensing modalities available on the physical robot. These sensors have included RGB cameras[44, 8, 1], depth cameras[44, 8], and LIDAR readings[44, 15]. In this work, we only employ RGB cameras for the following reasons. First, RGB cameras are cost-effective, power-efficient, and are passive sensors from an engineering perspective [24], which makes them more robust to environmental noise. Second, as we demonstrate in this work and as has been demonstrated in some past work [5, 6, 19], RGB sensors provide sufficient information regarding scene-level semantics for a suite of navigation and manipulation tasks. Third, the gap between simulation and the real world has shrunk for RGB observations due to improvements in physical cameras, far exceeding the quality provided by present-day real-world depth sensors.
Figure 4: Real Visual Observations: Our robot is shown on the Left and the observations from Nav Cam and Manip Cam are shown on the Top Right and Bottom Right respectively.
We use 2 RGB cameras mounted on the head of the robot to provide multi-view observation of the environment, as shown in Fig. 4. One camera referred to as the Navigation Camera (Nav Cam) faces the forward direction of the robot as it navigates around the world. The second camera referred to as the Manipulation Camera (Manip Cam) faces the arm of the Stretch robot. The Nav Cam predominantly provides sensory input for navigation but the Manip Cam can aid in manipulating the object but also in locating the target object during exploration. Each RGB observation has the shape of $384\times 224$. Even though the simulator[5, 20] provides photo-realistic RGB observations, there still exists a sizeable sim2real gap in visual observation. To counteract this gap, we apply augmentation over visual observation for robustness whose details can be found in the supplementary materials Sec 7.
@@
Proprioception: Since tasks in Daily Mobile Manipulation Task Suite demand manipulating objects in the environment, we feed the proprioception state of the robot to our controller as well. Specifically, the proprioception state is a 5-dimensional vector containing the lift of the arm base, the extension of the arm, and the rotation of the wrist.
4.3 HarmonicMM Architecture
Fig. 3 illustrates the network architecture of our proposed HarmonicMM controller. RGB observations from the two visual cameras are processed by a shared visual backbone followed by a separate small visual encoder consisting of two Conv Layers (visual encoder $E_{Nav}$ for visual observation from the navigation camera and visual encoder $E_{Manip}$ for visual observation from the manipulation camera) to produce visual embeddings. The visual embeddings are concatenated with the proprioceptive state of the robot. The resulting state embedding is then fed into a recurrent unit to add memory, which is subsequently followed by a linear layer to produce the next action, and details can be found in supplementary materials Sec 8.
Prior works have demonstrated that using a pre-trained visual backbone within the AI2-THOR environment is an effective approach for improving performance [19, 5]. Furthermore, employing a pre-trained visual backbone facilitates generalization to the real world. Past works [6] have underscored the effectiveness of using pre-trained visual backbones to capture task-specific semantic information from visual observations and to transfer effectively to the real world. Inspired by these prior efforts, we utilize a frozen DINOv2 [29] encoder, which is pre-trained on web-scale image data with self-supervision, to preprocess the RGB visual observations.
4.4 Reward Function
For all tasks in our Daily Mobile Manipulation Task Suite, the reward function is defined as follows:
where $w_{*}$ are coefficients for different reward terms. We provide high-level descriptions of each component of the function below and more details are provided in the ssupplementary materials Sec 10:
Navigation Reward $R_{\text{nav}}$: This reward is proportional to the distance the robot moved towards the target object (door/fridge for the Opening Door/Fridge task, or table for the Cleaning Table task). It provides a reach-target reward when the robot reaches the target and gets cut off afterward.
Manipulation Reward $R_{\text{manip}}$: This reward contains two sub-rewards. The first sub-reward encourages the robot to move its end-effector towards specific parts of the target object: the door knob for the Opening Door task, the center of the table for the Cleaning Table task, and the fridge handle for the Opening Fridge task. The second sub-reward is proportional to the progress the robot made towards task completion and offers a significant bonus upon successful task completion. Task progress is measured by the degree of openness of the door or fridge for Opening Door/Fridge tasks, and the percentage of dirt removed from the table for Cleaning Table tasks.
5 Experiment
[Opening Door (Pull) (From left to right)]Refer to caption
[Opening Door (Pull) (From left to right)]Refer to caption
[Opening Door (Push) (From left to right)]Refer to caption
[Cleaning Table (From left to right)]Refer to caption
Figure 5: Real World: We deployed the learned controller in a real apartment with a novel layout. Each row shows a single trajectory.
@@ 5 Experiment
Table 1: Task Completion Evaluation: We evaluate the task completion performance of different methods. Our method outperforms baselines on all tasks with higher Success Rates and higher Progresses.
@@
We evaluated our method on tasks in Daily Mobile Manipulation Task Suite as defined in Sec 3 both in simulation and in the real world. In the simulation, we simulated a Hello Robot Stretch [18] in diverse houses generated with ProcThor [5]. In the real world, we deployed the learned controller in a real apartment with multiple rooms without any adaptation or fine-tuning.
5.1 Training and Evaluation Details
We trained controllers for all methods for each task using DD-PPO [42, 34] for $10^{7}$ time steps using AllenAct[41]. Each experiment was repeated with $3$ different random seeds, and the mean and std over seeds were provided. We evaluated the policies for $300$ episodes per task per seed. Our evaluation were conducted in unseen houses with novel layouts. We used $\sim 50$, $\sim 200$, and $\sim 100$ unseen houses for Opening Door (Push and Pull), Cleaning Table, and Opening Fridge, respectively.
Robot Initialization Position in Scene: In mobile manipulation tasks, unlike tabletop manipulation where the robot base is stationary, the robot’s initial position significantly influences its exploration ability and overall performance. If the robot starts too close to the target object, it may struggle to adjust its pose adequately, hindering task completion. On the other hand, initializing the robot too far from the object might impede its ability to discover and learn specific behavior patterns effectively. In our study, we focus on mobile manipulation within everyday environments, ensuring that objects of interest are positioned within a practical range of the robot. In the task of Cleaning Table, the robot is initialized at a distance of $1\sim 4$ of the table. For the Opening Door/Fridge tasks, the robot begins within $1\sim 3$ from the door/fridge.
Metrics: We evaluate the performance of all methods with the following metrics:
  • Success Rate: For the tasks of opening doors and fridges, we considered an episode successful if the door or fridge was opened more than $90\%$ and $70\%$ respectively since fridges are normally located in a narrower space and human normally don’t fully open the fridge. For the task of cleaning the table, we considered the episode successful if the table was cleaned more than $75\%$.
  • Progress: This metric indicates how well the agent is making progress toward completing the task. Although a softer metric than the success rate, it reveals the agent’s progression towards task completion. Our tasks are relatively complex and require many steps to be completed, making this a particularly insightful metric, especially in cases where the agent does not reach a high completion rate but makes meaningful progress.
  • Progress Speed: This metric reflects the efficiency of the agent in making progress toward task completion. It is calculated by:
    Progress Speed=Task ProgressEpisode Length/Max Episode $\text{Progress Speed}=\frac{\text{Task Progress}}{\text{Episode Length}/\text{Max Episode Length}}$
  • Episode Length: This metric indicates the average number of steps the agent takes to complete the task. If the agent takes more than $500$ steps, we terminate the episode and consider that task as a failed example.
Table 2: Efficiency Evaluation: We evaluate the efficiency of different controller. Our method outperforms baselines on all tasks with higher Progress Speed and shorter Episode Length (Eps-Length).
@@ 5.1 Training and Evaluation Details
Table 3: Ablation Study: We ablate the necessity of proprioception information on Opening Door tasks. Proprioception information is necessary for tasks requiring precise placement of the end-effector and helpful for other tasks.

### v3

@@ Abstract
Recent advancements in robotics have enabled robots to navigate complex scenes or manipulate diverse objects independently. However, robots are still impotent in many household tasks requiring coordinated behaviors such as opening doors. The factorization of navigation and manipulation, while effective for some tasks, fails in scenarios requiring coordinated actions. To address this challenge, we introduce, HarmonicMM, an end-to-end learning method that optimizes both navigation and manipulation, showing notable improvement over existing techniques in everyday tasks. This approach is validated in simulated and real-world environments and adapts to novel unseen settings without additional tuning. Our contributions include a new benchmark for mobile manipulation and the successful deployment with only RGB visual observation in a real unseen apartment, demonstrating the potential for practical indoor robot deployment in daily life.
Figure 1: Harmonic Mobile Manipulation: In this work, we address diverse mobile manipulation tasks integral to human’s daily life. Trained in a photo-realistic simulation, Our controller effectively accomplishes tasks through harmonious mobile manipulation in a real-world apartment featuring a novel layout, without any fine-tuning or adaptation.
00footnotetext: Work done during internship at PRIOR @ Allen Institute for AI00footnotetext: More results on https://rchalyang.github.io/HarmonicMM
I Introduction
The field of robot learning has traditionally treated robot navigation and manipulation as separate domains, leading to significant advancements in each. On one hand, agents have been trained to efficiently explore and navigate environments ([1, 2, 3, 4]), while on the other, substantial progress has been made in performing complex manipulation tasks, such as handling articulated objects with static arms in tabletop settings ([5, 6, 7, 8]).
However, in household settings, many tasks require coordinated body movements while simultaneously manipulating objects with our arms. To address this challenge, recent efforts have focused on enhancing robotic agents with the ability for mobile manipulation. The most prominent approach is invoking navigation and manipulation as separate modules. This separation is achieved either by utilizing a high-level semantic controller, such as a large language model  ([9, 10]), or by estimating low-level subgoals using motion planners [11]. In these works, behaviors are split into distinct navigation and manipulation skills.
@@ I Introduction
  • Successful transfer of agents trained in simulation to real-world with only RGB visual observation.
@@
II Related Work
Embodied AI Benchmark. Over the past few years, a variety of standard benchmarks have emerged to assess progress in the field of embodied AI.
@@ II Related Work
These benchmarks primarily focus on high-level planning, scene understanding, and basic interaction tasks like object navigation and house rearrangement ([13, 12, 14, 15, 16, 17, 18, 19, 20]) and generally limit physical world interactions to straightforward pick-and-place actions. Deitke et al.[12] have created diverse houses for training embodied agents to navigate through various everyday scenes. Li et al.[21] tackles various household tasks using predefined motion primitives. In contrast to these benchmarks, our work extends beyond simple pick-and-place tasks. Our robot is designed to perform complex mobile manipulations, requiring tight coordination between navigation and manipulation.
Robotics Manipulation. In robotic manipulation, the focus has traditionally been on either fixed-base tabletop tasks or elementary mobile manipulations in controlled environments ([22, 23, 24, 25, 26, 27, 28, 29, 30]). Mu et al.[24] have developed benchmarks for straightforward mobile manipulation tasks in open spaces. Gu et al.[31] have demonstrated how robots can clean tables using imitation learning in a tabletop setup. While some works have explored basic mobile manipulations, such as door opening in lab settings [28], our efforts extend to more dynamic tasks requiring the robot to actively navigate and interact with its environment, pushing the boundaries of traditional mobile manipulation research.
@@
Mobile Manipulator. In the realm of mobile manipulation, previous studies have designed task-specific mobile manipulators using optimization techniques[32] or employing Task-and-motion-planning (TAMP) methods ([33, 34]). Lew et al.[32] developed a table-cleaning controller using whole-body trajectory optimization. While effective for their intended purposes, these methods tend to be narrowly focused with limited scalability and adaptability to different scenarios. Our pipeline could apply to a wide range of tasks, beyond the constraints of task-specific methodologies and offer a more flexible solution for various mobile manipulation challenges.
@@
Another significant line of work involves learning-based methods ([11, 35]), which offer better generalization in task completion. These methods diverge into two sub-categories: short-horizon whole-body control of robots ([36, 37, 38]), and long-horizon tasks solved at the scene level with predefined motion primitives in an iterative or two-stage manner ([39, 11, 40, 41, 9, 25, 26]). Yokoyama et al.[39], Xia et al.[11] and Ahn et al.[9] exemplify solving long-horizon mobile manipulation tasks using predefined navigation and manipulation skills, either iteratively or in a two-stage approach. Notably, our work proposes a unified learning framework that integrates navigation and manipulation in a seamless manner, addressing the limitations of predefined primitives by focusing on learning adaptable skills for complex, everyday tasks.
III Daily Mobile Manipulation Task Suite
@@
Figure 2: HarmonicMM Network Architecture (Left): Our HarmonicMM controller takes robot proprioception and multi-view visual observations as input and output navigation and manipulation commands at the same time. Real Visual Observations (Right): Our robot is shown on the Left and the observations from Nav Cam and Manip Cam are shown on the Top Right and Bottom Right respectively.
Figure 3: Our controller controls all DOF of the robot at every step.
Humans, on a daily basis, perform long-horizon tasks that necessitate close coordination between navigating the body and manipulating objects within the environment. While we can effortlessly coordinate actions between our hands and feet, teaching robots to do the same is a very challenging problem. To study this problem, first, we introduce a suite of tasks, named the Daily Mobile Manipulation Task Suite, consisting of atomic tasks that one might want a household robot to perform, as in Fig. 1. These tasks require an agent to navigate their world while manipulating objects.
1) Opening Door. In most living spaces, navigating through doors is a key task. This involves identifying and operating the door handle by pulling or pushing while adjusting the body position to avoid collision with the door. The suite includes both Opening Door (Push) and Opening Door (Pull).
2) Cleaning Table. A common chore is cleaning surfaces like tables and counters which requires coordinated movement to clean the entire surface area effectively. The suite features Cleaning Table to represent such activities.
3) Opening Fridge. Operating household appliances is crucial for a robotic assistant. This task involves locating and opening an appliance like a fridge, including navigation and manipulation. The suite includes Opening Fridge as an example of such tasks.
Training Environments. Our training leverages the ProcTHOR[12] environment, preparing agents for our task suite with an emphasis on bridging the simulation-to-reality (Sim2Real) gap through diverse and photorealistic simulations. For Opening Door tasks, we use 700 houses featuring two-room configurations with openable doors. Cleaning Table is trained with 2000 houses, varying from one to three rooms, each equipped with a table. For fridge-opening scenarios, 1400 houses with realistically furnished kitchens are employed. Environmental factors like textures and lighting are randomized to enhance generalization capabilities, ensuring effective real-world policy application.
IV Harmonic Mobile Manipulation (HarmonicMM)
Efficiently coordinating navigation and manipulation in mobile robots presents a significant challenge, particularly when using RL in large action spaces. Prior studies often split tasks into discrete navigation and manipulation stages, or alternately cycle between these modes, limiting the action space at any given timestep ([9, 26, 17, 11, 36, 42]). This formulation of the action space prevents the robot from performing challenging tasks such as those in Daily Mobile Manipulation Task Suite. Our approach employs the Stretch [43], which navigates and manipulates concurrently, leveraging a 5-dim continuous action space for integrated base and arm movements. As shown in Fig. 3, our action space includes base movements (translation and rotation), arm height and extension adjustments, and wrist rotation, with an additional dimension for grasp actions in specific tasks like door or fridge opening.
IV-A Simulated Environment.
@@ IV Harmonic Mobile Manipulation (HarmonicMM)
Our simulation utilizes ProcThor [12] for its superior visual diversity, addressing a gap in environments like IsaacGym [44] and Mujoco [45] that lack extensive visual variation critical for real-world skill transfer. To enable the robot to learn the diverse tasks described in Sec III, We augmented the simulation with procedural door/fridge opening logic and photo-realistic table cleaning functionality.
@@
Procedural Door/Fridge Opening: One of our contributions is modifying ProcThor to enable physically realistic articulation manipulation. In the ProcThor[12] simulation environment, doors and fridges are treated as semantic objects, and their open and closed states are not governed by any physical rules. Instead, we estimate their status at each simulation step based on the robot’s actions. For pushing actions, the system calculates the direction the robot’s end-effector is moving towards. If this direction aligns with the door’s normal vector, the door is opened incrementally in the same direction as the end-effector’s movement. In the case of pulling actions, the door’s movement is determined after the robot command is executed. The door opens in small segments if the end-effector’s movement direction aligns with the door’s normal vector, provided the robot has already secured the doorknob with a magnetic grasper. Additionally, collision detection algorithms are in place to prevent the door from colliding with the robot’s body during these interactions.
@@ IV-B Robot Perception
We equip the robot with both RGB visual observations as well as the proprioceptive state of the robot, which is available on most robots and easy to obtain.
Multi-camera RGB Observations: In contrast to prior work utilizing a mix of RGB, depth cameras, and LIDAR for mobile manipulation, our approach exclusively uses RGB cameras due to their cost-efficiency, power efficiency, and robustness against environmental noise. RGB cameras have been proven to provide ample scene semantics for navigation and manipulation tasks, supported by advances that have narrowed the simulation-to-reality visual gap ([11, 38, 9, 46, 12, 2, 1]). Our robot is equipped with two RGB cameras: a Navigation Camera (Nav Cam) for forward observation and a Manipulation Camera (Manip Cam) aimed at the robot’s arm, enabling multi-view environmental awareness, as shown in Fig.  2. Despite the photorealism in simulation ([12, 13]), we address the remaining Sim2Real visual discrepancies through augmentation, further detailed in the appendix (Sec VII).
@@
Proprioception: For tasks requiring object manipulation in Daily Mobile Manipulation Task Suite, our controller utilizes a 5-dimensional proprioception vector, detailing the arm base’s lift, arm extension, and wrist rotation, to enhance manipulation capabilities.
@@
TABLE I: Task Completion Evaluation (Top): Our method outperforms baselines with higher Success Rates and higher Progress. Efficiency Evaluation (Bottom): Our method outperforms baselines with higher Progress Speed and shorter Episode Length (Eps-Length).
@@
IV-C HarmonicMM Architecture
As shown in Fig. 2, HarmonicMM starts with a DINOv2 visual backbone that processes input from both RGB cameras. These inputs are further refined by two convolutional visual encoders, $E_{Nav}$ and $E_{Manip}$, for navigation and manipulation visual observations, respectively, to generate specific visual embeddings. Visual embeddings are concatenated with the robot’s proprioception and then fed into a recurrent unit to add memory, followed by a linear layer to produce the robot action. More details can be found in Sec VIII. Leveraging a pre-trained visual backbone, as demonstrated by previous research within AI2-THOR ([1, 12, 2, 47]), offers significant advantages in performance and real-world generalization. This choice is informed by the encoder’s proven ability to capture detailed semantic information from visual observations and its success in bridging the simulation-to-reality gap, thanks to its self-supervised training on a diverse web-scale dataset.
IV-D Reward Function
We use the following reward function for all tasks:
where $w_{*}$ are coefficients for different reward terms. We provide high-level descriptions of each term below and more details are provided in Sec X.
Navigation Reward $R_{\text{nav}}$: This reward is proportional to the distance the robot moved towards the target object (door/fridge for the Opening Door/Fridge, or table for Cleaning Table). It provides a reach-target reward when the robot reaches the target and gets cut off afterward.
Manipulation Reward $R_{\text{manip}}$: This reward contains two sub-rewards. The first sub-reward encourages the robot to move its end-effector towards the doorknob, the center of the table, and the fridge handle for Opening Door, Cleaning Table, and Opening Fridge tasks respectively. The second sub-reward is proportional to the progress the robot made towards task completion and offers a significant bonus upon successful task completion. Task progress is measured by the degree of openness of the door or fridge for Opening Door/Fridge tasks, and the percentage of dirt removed from the table for Cleaning Table tasks.
V Experiment
Figure 4: Real World: We deployed the learned controller in a real apartment with a novel layout. Each row shows a single trajectory (from left to right) corresponding to Opening Door (Pull), Opening Door (Pull), Opening Door (Push), and Cleaning Table, respectively.
We evaluated our method on Daily Mobile Manipulation Task Suite both in simulation and in the real world. We deployed the policy (learned in simulation) in a real apartment with multiple rooms without any adaptation or fine-tuning.
V-A Training and Evaluation Details
We trained the policies for all methods for each task using DD-PPO( [48]) for $10^{7}$ steps using AllenAct[49]. Each experiment was repeated with $3$ different seeds. We evaluated the policies on $1200$ episodes ($300$ per task). Our evaluations were conducted in 350 unseen houses with novel layouts.
Robot Initialization Position in Scene: Recognizing the impact of the initial position of the robot on mobile manipulation, we optimized the robot’s start points to improve exploration efficiency and task performance. If the robot is too close to the object, it struggles to adjust its pose, while, initializing the robot too far from the object negatively impacts its ability to learn required skills. We place the robot $1-4$ meters away from the table for Cleaning Table, and $1-3$ meters away from the door/fridge for Opening Door/Fridge tasks.
Metrics: We evaluate the performance of all methods with:
Progress: This metric (as defined in IV-D) indicates how much progress the agent made toward completing the task. Our tasks are complex and long-horizon, making this an insightful metric, especially where the agent does not reach a high completion rate but makes meaningful progress.
Success Rate: Success was defined by the task progress: over $90\%$ for door opening, $70\%$ for fridge opening (since normally fridges are located in a narrower space and humans don’t fully open the fridge), and $75\%$ for table cleaning.
Episode Length: Reflects the average steps required for task completion, with episode length capped at $500$ steps.
Progress Speed: This metric reflects the efficiency of the agent in making progress toward task completion. It is calculated by: Task ProgressEpisode Length/Max Episode $\frac{\text{Task Progress}}{\text{Episode Length}/\text{Max Episode Length}}$
TABLE II: Proprioception Ablation: We ablate the necessity of proprioception information on Opening Door tasks. Proprioception information is necessary for tasks requiring precise placement of the end-effector and helpful for other tasks.